# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v2

@@ C.15 What the bounds do not include
The surrogate and budgeted information bounds describe model limits: they do not include (i) finite-sample dispersion control from the sample cap (we encode it population-wise via $\rho$) or (ii) the oracle first-stage uncertainty (added separately by OUA).
C.16 Citations for technical ingredients
Semiparametric efficiency and one-step/TMLE: Bickel et al. (1993); van der Vaart & Wellner (2000); van der Laan & Rose (2011); Tsiatis (2006).
Isotonic regression / PAVA and order-restricted inference: Ayer et al. (1955); Barlow et al. (1972); Robertson et al. (1988); Banerjee (2001).
Majorization theory: Hardy et al. (1952); Marshall et al. (2011).
Calibration for OPE: Kallus & Mao (2020); van der Laan et al. (2025; 2024).
Jackknife/bootstraps for dependence: Künsch (1989); Politis & Romano (1994).
Projections and non-expansive maps: Bauschke & Combettes (2017).
@@ D.2 Judge calibration, coverage, and drift
Partition $S$ into $B$ bins; for bin $b$, plot the bin mean of $R=\hat{f}(Z(S))$ against the oracle mean of $Y$, with 95% binomial intervals. Report a Brier-style reliability term and the OOF RMSE, with a 1-SE model-selection overlay (monotone vs. two-stage).
@@ D.3 DR orthogonality and decomposition
D.4 Uncertainty: IF variance and OUA addition
@@
with $\widehat{\operatorname{Var}}_{\mathrm{cal}}$ from the oracle jackknife (App. A.7). Report the OUA share $\widehat{\operatorname{Var}}_{\mathrm{cal}}/\widehat{\operatorname{Var}}_{\mathrm{total}}$ and, optionally, a dependence-robust alternative (below).
@@ D.6 Policy-wise mean transport diagnostic
By 3, $H_{0,\pi^{\prime}}$ holds if and only if the surrogate-estimated policy value equals the oracle value: $\mathbb{E}_{\pi^{\prime}}[f(S,X)]=\mathbb{E}_{\pi^{\prime}}[Y]$. Rejecting $H_{0,\pi^{\prime}}$ implies that surrogate-only evaluation is systematically biased for policy $\pi^{\prime}$: the calibration does not transport.
OUA-corrected inference.
To account for uncertainty in estimating $f$, the confidence interval for the mean residual should incorporate OUA variance (Section 3.5). In practice, we refit $f^{(-k)}$ omitting each oracle fold and compute residuals under each refit, propagating calibration uncertainty into the test statistic.
@@ D.7 Gates: thresholds and actions
Mean sufficiency All Reliability diagram; regional residuals centered at zero Two-stage fallback with covariates; refresh oracle slice
Monotonicity in $S$ AutoCal-R OOF model selection: monotone vs. two-stage; 1-SE rule Use two-stage with response length covariate
@@
Transportability Direct Residual test: $\mathbb{E}[Y-f(S,X)]=0$ per policy (Bonferroni) Recalibrate per policy; collect policy-specific oracle data
@@ D.8 Planner: MDE and label/log budgets
We tabulate $\widehat{\mathrm{SE}}$ versus (n,m/n)(n,\,m/n) (labels per log) using Stacked-DR with OUA and annotate “iso-cost” lines for the label budget.
D.9 Reporting ledger (per policy/cohort)
Persist: (i) calibrator mode, OOF risk by tertiles, knots/levels (hash); (ii) SIMCal–W maps, stacking weights $\hat{\beta}$, guard $\rho$ and blend $\alpha$; (iii) ESS fraction, max-weight share, Hill index band, $A_{B}$; (iv) DR orthogonality score and CI; DM–IPS split; (v) OUA trace $\{\widehat{V}^{(-k)}\}$ and variance breakdown; (vi) filter counts (e.g., TF gaps) and an inclusion manifest of $x$-ids; (vii) multiplicity control (family, $q$, adjusted $p$).
D.10 Visualization primitives (for reproducible panels)
  • ESS/tails strip: bars for ESS fraction (baseline vs. SIMCal–W); dot for max-weight share; Hill band.
@@ D.10 Visualization primitives (for reproducible panels)
  • Uncertainty ring: pie of $\widehat{\operatorname{Var}}_{\mathrm{cal}}/\widehat{\operatorname{Var}}_{\mathrm{total}}$ (OUA share).
D.11 Optional: dependence–robust implementation details
@@
D.12 Compact gate pseudo-logic
@@
1: Compute weight/tail metrics: ESS fraction, max-share, Hill band; compute $A_{B}$; judge reliability/coverage; orthogonality score; OUA share
@@ E.3 Folds and cross–fitting
We use K=5K{=}5 folds by default. The fold map $F(i)$ is a stable hash of x_id modulo $K$, ensuring that all modules (AutoCal–R, SIMCal–W, DR nuisances) share identical OOF boundaries. Oracle folds are derived by intersecting $F(i)$ with $L_{i}{=}1$. The hash rule is $F(x\_id){=}\mathrm{hash}(x\_id)\bmod 5$.
@@ E.5 Persisted artifacts (per policy/cohort)
  • OUA: $\{\widehat{V}^{(-k)}\}_{k=1}^{K}$, $\widehat{\operatorname{Var}}_{\mathrm{cal}}$, $\widehat{\operatorname{Var}}_{\mathrm{total}}$.
@@ E.6 Reference run order (pseudocode)
build_tf_cache --policies <list> --dataset logs.parquet --out tf_cache.parquet
@@
# 2) SIMCal-W per policy (OOF project->stack->cap->translate-to-mean-one)
simcal_w --tf-cache tf_cache.parquet --scores S.parquet --rho 1.0 --folds 5 \
         --out weights.parquet
# 3) Estimation + IFs (Cal-IPS / DR-CPO)
estimate --rewards rewards.parquet --weights weights.parquet --folds 5 \
         --out estimates.parquet
# 4) IF-Stack (optional)
@@
# 5) OUA jackknife
oua --oracle-folds 5 --pipeline-config cfg.yaml --out variance.parquet
@@ E.7 Compute and resource model
  • SIMCal–W. $O(n\log n)$ for sort ++ $O(n)$ for PAVA per fold; covariance/stacking are tiny ($3{\times}3$).
  • DR-CPO. If $\hat{g}_{\pi^{\prime}}$ uses one rollout per $(X,\pi^{\prime})$, add $O(|\Pi^{\prime}|\,n\,\bar{T}^{\prime})$ tokens once; a light smoother amortizes Monte Carlo noise.
  • OUA. $K$ refits of AutoCal–R and re–runs of the pipeline; cache features to avoid recomputation.
@@ Appendix F Efficient Budget Allocation
Note: This first-order approximation assumes additive variance components with $O(1/m)$ calibration variance. While isotonic regression has nonparametric rates for function estimation, the mean functional $\mathbb{E}[f(S)]$ typically achieves $\sqrt{m}$ rates under smoothness. In practice, estimate $\sigma^{2}_{\mathrm{cal}}\approx m\cdot\widehat{\operatorname{Var}}_{\mathrm{cal}}$ and $\sigma^{2}_{\mathrm{eval}}\approx n\cdot\widehat{\operatorname{Var}}_{\mathrm{eval}}$ from the OUA decomposition, then apply the formula.
@@
The OUA Marginal Utility Diagnostic.
The optimality condition implies that resources are allocated efficiently when the marginal variance reduction per dollar is equal across both channels. We can express this in terms of the observed OUA Share $\omega=\operatorname{Var}_{\mathrm{cal}}/\operatorname{Var}_{\mathrm{total}}$. At the optimum, the Spend-Balance Rule holds:
@@
Using costs from our experiments: the gpt-4.1-nano-2025-04-14 judge is approximately $16\times$ cheaper than the gpt-5-2025-08-07 oracle, giving a cost ratio $c_{S}/c_{Y}\approx 0.064$. At $n=1000$ prompts with $m=50$ oracle labels (5% oracle fraction), the observed OUA share is $\omega\approx 55\%$.
@@
$\frac{\sigma^{2}_{\mathrm{cal}}}{\sigma^{2}_{\mathrm{eval}}}=\frac{\omega\cdot m}{(1-\omega)\cdot n}=\frac{0.55\times 50}{0.45\times 1000}=0.061$
@@
$\frac{m^{*}}{n^{*}}=\sqrt{0.064}\times\sqrt{0.061}\approx 0.25\times 0.25=6.2\%$
The empirical 5% oracle allocation is close to the variance-optimal 6.2%, confirming that the study’s cost-effective design did not substantially sacrifice precision.

### v3

@@ C.17 What the bounds do not include
The surrogate and budgeted information bounds describe model limits: they do not include (i) finite-sample dispersion control from the sample cap (we encode it population-wise via $\rho$) or (ii) the oracle first-stage uncertainty (added separately by calibration-aware inference).
C.18 Budgeted bound and IF stacking (OPE-specific results)
The following results apply to off-policy evaluation (IPS and DR estimators). Direct mode does not use importance weights and is unaffected by these bounds.
Budgeted bound (variance cap).
For $\rho\geq 1$, define
Theorem 4 (Budgeted information bound).
The optimal asymptotic variance under the cap $\rho$ equals $\operatorname{Var}(\phi^{(\rho)})$, which is nonincreasing in $\rho$ and satisfies $\lim_{\rho\to\infty}\operatorname{Var}(\phi^{(\rho)})=\operatorname{Var}(\phi_{\mathrm{sur}})$. If weight stabilization converges to $m^{\star}_{\rho}$ (with the guard), then calibrated DR attains $\operatorname{Var}(\phi^{(\rho)})$.
Theorem 5 (Influence-function stacking).
@@ D.2 Judge calibration, coverage, and drift
Partition $S$ into $B$ bins; for bin $b$, plot the bin mean of $R=\hat{f}(Z(S))$ against the oracle mean of $Y$, with 95% binomial intervals. Report a Brier-style reliability term and the OOF RMSE.
@@ D.3 DR orthogonality and decomposition
D.4 Uncertainty: IF variance and calibration uncertainty addition
@@
with $\widehat{\operatorname{Var}}_{\mathrm{cal}}$ from the oracle jackknife (App. A.7). Report the calibration uncertainty share $\widehat{\operatorname{Var}}_{\mathrm{cal}}/\widehat{\operatorname{Var}}_{\mathrm{total}}$ and, optionally, a dependence-robust alternative (below).
@@ D.6 Policy-wise mean transport diagnostic
By 4, $H_{0,\pi^{\prime}}$ holds if and only if the surrogate-estimated policy value equals the oracle value: $\mathbb{E}_{\pi^{\prime}}[f(S,X)]=\mathbb{E}_{\pi^{\prime}}[Y]$. Rejecting $H_{0,\pi^{\prime}}$ implies that surrogate-only evaluation is systematically biased for policy $\pi^{\prime}$: the calibration does not transport.
Calibration-uncertainty-corrected inference.
To account for uncertainty in estimating $f$, the confidence interval for the mean residual should incorporate calibration uncertainty (Section 3.5). In practice, we refit $f^{(-k)}$ omitting each oracle fold and compute residuals under each refit, propagating calibration uncertainty into the test statistic.
@@ D.7 Gates: thresholds and actions
Transport (T1) All (for valid levels) Mean residual test: $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$ per policy Fundamental requirement; test with audit; recalibrate or refuse levels if fail
Mean sufficiency (J2-M) Reward calibration Reliability diagram; regional residuals centered at zero Structural, sufficient for transport; two-stage fallback; not required if transport passes
Monotonicity in $S$ Reward calibration OOF RMSE comparison Default to two-stage; use monotone-only if covariates unavailable
@@
Assumption hierarchy. Transport (T1) is the fundamental requirement for valid surrogate-only level claims. Mean sufficiency (J2-M) and Prentice-style surrogacy are sufficient conditions that imply transport, but transport can hold without them—errors may cancel in expectation. With an audit slice, test transport directly; structural assumptions become optional. See Section C.15 for details.
@@ D.8 Planner: MDE and label/log budgets
We tabulate $\widehat{\mathrm{SE}}$ versus (n,m/n)(n,\,m/n) (labels per log) using Stacked-DR with calibration-aware variance and annotate ”iso-cost” lines for the label budget.
D.9 Weight Stabilization Visualization
Figure 10 illustrates the effect of weight stabilization on importance weight distributions across target policies.
[…]
D.10 Reporting ledger (per policy/cohort)
Persist: (i) calibrator mode, OOF risk by tertiles, knots/levels (hash); (ii) weight stabilization maps, stacking weights $\hat{\beta}$, guard $\rho$ and blend $\alpha$; (iii) ESS fraction, max-weight share, Hill index band, $A_{B}$; (iv) DR orthogonality score and CI; DM–IPS split; (v) calibration uncertainty trace $\{\widehat{V}^{(-k)}\}$ and variance breakdown; (vi) filter counts (e.g., TF gaps) and an inclusion manifest of $x$-ids; (vii) multiplicity control (family, $q$, adjusted $p$).
D.11 Visualization primitives (for reproducible panels)
  • ESS/tails strip: bars for ESS fraction (baseline vs. weight stabilization); dot for max-weight share; Hill band.
@@ D.11 Visualization primitives (for reproducible panels)
  • Uncertainty ring: pie of $\widehat{\operatorname{Var}}_{\mathrm{cal}}/\widehat{\operatorname{Var}}_{\mathrm{total}}$ (calibration uncertainty share).
D.12 Optional: dependence–robust implementation details
@@
D.13 Compact gate pseudo-logic
@@
1: Compute weight/tail metrics: ESS fraction, max-share, Hill band; compute $A_{B}$; judge reliability/coverage; orthogonality score; calibration uncertainty share
@@ E.3 Folds and cross–fitting
We use K=5K{=}5 folds by default. The fold map $F(i)$ is a stable hash of x_id modulo $K$, ensuring that all modules (reward calibration, weight stabilization, DR nuisances) share identical OOF boundaries. Oracle folds are derived by intersecting $F(i)$ with $L_{i}{=}1$. The hash rule is $F(x\_id){=}\mathrm{hash}(x\_id)\bmod 5$.
@@ E.5 Persisted artifacts (per policy/cohort)
  • Calibration-aware inference: $\{\widehat{V}^{(-k)}\}_{k=1}^{K}$, $\widehat{\operatorname{Var}}_{\mathrm{cal}}$, $\widehat{\operatorname{Var}}_{\mathrm{total}}$.
@@ E.6 Reference run order (pseudocode)
build_tf_cache --policies <list> --dataset logs.parquet \
               --out tf_cache.parquet
@@
# 2) Weight stabilization per policy
#    (OOF project -> stack -> cap -> translate-to-mean-one)
weight_stabilize --tf-cache tf_cache.parquet --scores S.parquet \
                 --rho 1.0 --folds 5 --out weights.parquet
# 3) Estimation + IFs (Calibrated-IPS / Calibrated-DR)
estimate --rewards rewards.parquet --weights weights.parquet \
         --folds 5 --out estimates.parquet
# 4) Influence-function stacking (optional)
@@
# 5) Calibration-aware jackknife
cal_jackknife --oracle-folds 5 --pipeline-config cfg.yaml \
              --out variance.parquet
@@ E.7 Compute and resource model
  • Weight stabilization. $O(n\log n)$ for sort ++ $O(n)$ for PAVA per fold; covariance/stacking are tiny ($3{\times}3$).
  • Calibrated DR. If $\hat{g}_{\pi^{\prime}}$ uses one rollout per $(X,\pi^{\prime})$, add $O(|\Pi^{\prime}|\,n\,\bar{T}^{\prime})$ tokens once; a light smoother amortizes Monte Carlo noise.
  • Calibration-aware inference. $K$ refits of reward calibration and re–runs of the pipeline; cache features to avoid recomputation.
@@ Appendix F Efficient Budget Allocation
Note: This first-order approximation assumes additive variance components with $O(1/m)$ calibration variance. While isotonic regression has nonparametric rates for function estimation, the mean functional $\mathbb{E}[f(S)]$ typically achieves $\sqrt{m}$ rates under smoothness. In practice, estimate $\sigma^{2}_{\mathrm{cal}}\approx m\cdot\widehat{\operatorname{Var}}_{\mathrm{cal}}$ and $\sigma^{2}_{\mathrm{eval}}\approx n\cdot\widehat{\operatorname{Var}}_{\mathrm{eval}}$ from the calibration-aware variance decomposition, then apply the formula.
@@
The Calibration Uncertainty Marginal Utility Diagnostic.
The optimality condition implies that resources are allocated efficiently when the marginal variance reduction per dollar is equal across both channels. We can express this in terms of the observed calibration uncertainty share $\omega=\operatorname{Var}_{\mathrm{cal}}/\operatorname{Var}_{\mathrm{total}}$. At the optimum, the Spend-Balance Rule holds:
@@
Using costs from our experiments: the gpt-4.1-nano-2025-04-14 judge is approximately $16\times$ cheaper than the gpt-5-2025-08-07 oracle, giving a cost ratio $c_{S}/c_{Y}\approx 0.064$. At $n=1000$ prompts with $m=50$ oracle labels (5% oracle fraction), the observed calibration uncertainty share is $\omega\approx 90\%$.
@@
$\frac{\sigma^{2}_{\mathrm{cal}}}{\sigma^{2}_{\mathrm{eval}}}=\frac{\omega\cdot m}{(1-\omega)\cdot n}=\frac{0.90\times 50}{0.10\times 1000}=0.45$
@@
$\frac{m^{*}}{n^{*}}=\sqrt{0.064}\times\sqrt{0.45}\approx 0.25\times 0.67=17\%$
[…]
@@
Appendix G Code Generation Experiment (MBPP)
To validate that CJE generalizes beyond preference judgments, we evaluated on MBPP code generation (Austin et al., 2021), where ground truth is objective: unit tests either pass or fail.
Setup.
We use 257 MBPP problems (974 samples across 5 epochs) with five policies: base (GPT-4o standard, 87.4% pass rate, used for calibration), enhanced (GPT-4o with expert prompting, 87.2%), terse (GPT-4o with minimal prompting, 87.4%), mini (GPT-4o-mini, 80.9%), and unhelpful (deliberately buggy, 56.0%). The judge is GPT-4o-mini predicting pass probability (0–100, normalized to 0–1). Oracle labels are actual unit test results. Unlike Arena (where response length confounds judge scores), MBPP has no strong covariate effects, so we use monotone-only calibration.
Transport validation.
We apply the mean transport test (4) to each target policy. Two policies fail:
  • mini: Mean residual 0.065-0.065 (p=0.006p{=}0.006). The judge model (GPT-4o-mini) is identical to this policy, creating a judge–policy confound that violates transport.
  • unhelpful: Mean residual 0.280-0.280 (p<0.001p{<}0.001). Low-support saturation: 31% of samples score below base’s 10th percentile, where calibrator training density is sparse.
Both policies are excluded from RMSE computation; including them would bias results.
Results.
Table 12 shows CJE performance across oracle fractions for the two transport-valid policies (enhanced, terse) at full sample size (n=974n{=}974). Calibration provides consistent RMSE reduction across all oracle fractions: even at 5% oracle (${\sim}50$ labels), CJE achieves 18% improvement over naive estimation (raw judge scores without calibration). At 10%+ oracle, improvements reach 32–84%.
Table 12: CJE on MBPP code generation: oracle fraction tradeoff. RMSE computed over two transport-valid policies (enhanced, terse) at n=974n{=}974. Unlike Arena where 5% oracle suffices for ranking, MBPP shows diminishing returns: 10% oracle captures most of the benefit.
Oracle % # Labels CJE Naive (uncalib.) Improvement
5% 49 0.039 0.048 +18%
10% 97 0.033 0.048 +32%
25% 244 0.019 0.048 +60%
50% 487 0.014 0.048 +70%
100% 974 0.008 0.048 +84%
Comparison to Arena.
Both experiments validate the oracle fraction tradeoff, but with different thresholds:
  • Arena: 5% oracle achieves 94% ranking accuracy; calibration benefit plateaus early.
  • MBPP: 5% oracle provides 18% RMSE reduction; 10% provides 32%; benefit continues scaling to 84% at full oracle.
The difference likely reflects task structure: Arena evaluates preferences (ordinal, high variance) while MBPP evaluates correctness (binary, low variance). Binary oracles require more labels to stabilize calibration.
Judge–policy confound.
The mini transport failure highlights a methodological risk: when a model judges its own outputs, self-preference or style-alignment biases can corrupt calibration. Practitioners should use external judges or validate transport explicitly.
Takeaway.
[…]